### v1

# Identifying and Improving Disability Bias in GAI-Based Resume Screening

Abstract.
As Generative AI rises in adoption, its use has expanded to include domains such as hiring and recruiting. However, without examining the potential of bias, this may negatively impact marginalized populations, including people with disabilities. To address this important concern, we present a resume audit study, in which we ask ChatGPT (specifically, GPT-4) to rank a resume against the same resume enhanced with an additional leadership award, scholarship, panel presentation, and membership that are disability related. We find that GPT-4 exhibits prejudice towards these enhanced CVs. Further, we show that this prejudice can be quantifiably reduced by training a custom GPTs on principles of DEI and disability justice. Our study also includes a unique qualitative analysis of the types of direct and indirect ableism GPT-4 uses to justify its biased decisions and suggest directions for additional bias mitigation work. Additionally, since these justifications are presumably drawn from training data containing real-world biased statements made by humans, our analysis suggests additional avenues for understanding and addressing human bias.
Resume Audit, Bias, Ableism
copyright: nonejournalyear: 2024doi: conference: ; 2024; isbn: ccs: Social and professional topics People with disabilitiesccs: Social and professional topics Employment issuesccs: Computing methodologies Artificial intelligence
1. Introduction
Generative Artificial Intelligence (GAI) is being increasingly used for workforce recruiting and human resource management (e.g., (Lee, 2023; Nemirovsky et al., 2021; Rane, 2023; Rathnayake and Gunawardana, 2023; Rane, 2023)). One common example is resume screening, where artificial intelligence is used to rank resumes, a task for which large language models (LLMs) such as ChatGPT are increasingly used (e.g., (Lever Team, 2023; rec, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023)). GAI’s advantages include optimizing the potentially time-consuming process of screening resumes to a fraction of what a purely human-driven review process would take (Tran et al., 2023), and accurately summarizing lengthy application materials to highlight a candidate’s strengths and weaknesses (Lee, 2023). However, there is a danger in using AI for hiring: AI-based resume filtering and recruitment systems are biased, for example against candidates of diverse genders (Gagandeep et al., 2023; Kodiyan, 2019) and ages (Harris, 2023). Prior work has also highlighted potential risks of disability bias in hiring (e.g., (Buyl et al., 2022; Tilmes, 2022)). However, no prior work has quantified the amount of bias due to disability in deployed AI systems.
Further, no work we are aware of has demonstrated a way to reduce disability bias in resume screening. While human involvement and collaboration has been posed as a solution to general AI-created bias (Van den Broek et al., 2021), existing research shows that even experienced recruiters with expressed skepticism for AI-based solutions may default to accepting AI-based feedback when receiving inconsistent recommendations from a system (Lacroux and Martin-Lacroux, 2022). This highlights the importance of addressing bias in the AI systems themselves, in addition to any human interventions.
This article addresses these gaps by quantifying, and then reducing, bias in AI-based resume screening specifically for people with disabilities. Disabled people, of whom there are 42.5 million in the United States (Leppert and Schaeffer, 2023), already face significant barriers to employment, including fewer callbacks and inequality in the labor market (Ameri et al., 2015; Bjørnshagen and Ugreninov, 2021). Any ableist(Cherney, 2011) bias in AI-based hiring systems could exacerbate such employment barriers. Yet these systems are already in use (Lever Team, 2023; rec, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023). It is urgent that we understand, evaluate, and mitigate bias present in GAI-based resume screening for people with disabilities. Therefore, this work seeks to address the following research questions:
RQ1: DisabilityDifference:
Does a GAI-based resume screening exhibit bias against resumes that mention disability compared to those that do not? Is this bias different depending on the type of disability mentioned in the resume?
RQ2: BiasReduction:
Does a GAI trained on DEI principals exhibit reduced bias in comparison to a generic GAI?
RQ3: BiasExplanation:
Do GAI explanations of rankings provide evidence for potential sources or types of bias?
After summarizing related work on bias in AI-based hiring, and quantification of bias in hiring through resume audits in Related Work (Section 2), we describe our mixed-methods resume audit study method in Methods (Section 3). We test for bias by asking ChatGPT, and a DEI and disability-justice (Berne et al., 2018) trained custom GPT, to complete a series of ranking tasks comparing a control resume to a resume enhanced with a disability-related leadership award, scholarship, panel presentation and organizational membership. We vary the type of disability mentioned in the enhanced resume, and ask both GAIs to rank each control/enhanced pair 10 times, providing an explanation each time. Our findings quantitatively demonstrate statistically significant bias in ranking, significant differences in amount of bias across different disabilities, and that training can significantly reduce bias (Section 4). Further, our qualitative analysis of GPT’s explanations for its rankings (Section 4.4) demonstrates both direct and indirect ableist reasoning.
To summarize, our work takes a systematic approach to quantify ableist bias and complement these findings with qualitative evidence. Our study method is novel because of its use of direct comparison, since our use of GPT allows ranking a standard resume against a resume enhanced with disability-related items, instead of simply measuring callbacks or responses to a single resume. Our results are especially important because we demonstrates bias using AI tools that are currently being used to rank resumes. Further, our study is the first resume audit study to uncover the reasoning behind such biases, since our use of GPT also supports collection and analysis of the rationale behind the rankings. Since “…society’s racism, misogyny, ableism, etc., tend to be overrepresented in training data …[an LLM] that has been trained on such data will pick up these kinds of problematic associations” (Bender et al., 2021). Thus, our qualitative analysis of GAI-produced text may help to uncover biased reasoning that also impacts human judgment, something rarely part of resume audit studies. Based on these findings, we highlight important avenues for further work in Discussion and Recommendations (Section 5) and Ethical Guidance (Section 7). Our recommendations could pave the path for future efforts to mitigate bias and make GAI-based recruiting systems truly useful in equitable hiring.
2. Related Work
Hiring bias is an unfortunate reality, and has been linked to the unconscious and sometimes conscious mental process which influences the evaluation of candidates (Harvard Business School, 2020), including biases based on factors such as gender, race, and ethnicity (e.g., (Kessler et al., 2019; Ock, 2022; Thanasombat and Trasviña, 2005; Zarb, 2022)). Studies have shown that aspects of a hiring profile, such as the applicant’s name, can indicate an applicant’s ethnicity and trigger a biased response (Thanasombat and Trasviña, 2005). From the perspective of a disabled job-seeker, bias is an unfortunate reality that is sometimes mitigated by controlling when, whether and how they disclose their disabilities (e.g., (Alexandrin et al., 2008; Marshall et al., 2020; Evans, 2019; Lyons et al., 2018; Grimes et al., 2019; Charmaz, 2010; Kulkarni, 2022)). Some studies propose addressing bias by raising awareness of unconscious bias and its implications, or implementing methods to classify candidates in a way that minimizes the impact of bias (Harvard Business School, 2020; Bohnet et al., 2020).
The use of AI in hiring has many potential benefits, from actionable and constructive feedback for job seekers (Nemirovsky et al., 2021; Du et al., 2023) to time savings for recruiters (Tran et al., 2023). Generative AI in particular can quickly summarize and highlight important aspects of applications (Tran et al., 2023; Lee, 2023). However, AI has been demonstrated to reproduce human biases, spurring a movement in some countries such as the EU AI Act, which monitors its use for critical areas such as employment (Parliament, 2023). In the U.S., widely available AI systems are regularly used in candidate tracking with limited oversight. The emergence of low-cost GAI tools such as ChatGPT, which are in use today recruiting and hiring (e.g., (Lever Team, 2023; rec, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023)), has created an urgent need to understand the ethics and risks of such systems. Such risks have been documented in systems pre-dating and post-dating the advent of readily available generative AI, and we highlight some of the pertinent concerns in Section 2.1, showing that very little is known about disability bias in this context. One well understood way of quantifying bias is a resume audit study. Such studies have been traditionally used to measure human bias (e.g., biases due to racial identity (Bertrand and Mullainathan, 2004; Kang et al., 2016; Oreopoulos and Dechief, 2012), degree of ethnic identification (Derous and Ryan, 2012), queer identity and participation in LGBTQ+ organizations (Mishel, 2016), and disability status (l’Horty et al., 2022; Lippens et al., 2023)) but have also been used to measure AI bias in resume screening (Rhea et al., 2022). Section 2.2 introduces the method and summarizes some relevant findings, highlighting that although disability has been studied in a resume audit (Ameri et al., 2015), this has not yet translated into resume audits of AI (Rhea et al., 2022).
2.1. AI/ML Hiring Tools and Bias
Even before the wide-spread adoption of generative AI, AI and machine learning were widely employed in hiring, and widely studied due to concerns about bias (Lacroux and Martin-Lacroux, 2022; Pena et al., 2020; Deshpande et al., 2020; Mujtaba and Mahapatra, 2019). These biases are thought to exist because the datasets for the models carry human biases themselves (Deshpande et al., 2020). Existing research explores the different dimensions of bias present in these AI models and the tools that use them, such as biases based on socio-linguistic ability, age, gender, and race (Deshpande et al., 2020; Harris, 2023; Vaishampayan et al., 2023). It was also found that the AI/ML models could discern characteristics of a person from their resume when details weren’t explicitly given (Pena et al., 2020). This has led to efforts in reducing the bias by these algorithms through masking characteristics in resumes or creating a more human-centered AI algorithm or tool design (Vaishampayan et al., 2023; Gagandeep et al., 2023).
With the advent of generative AI and its pervasive issues with bias and ethical concerns (McDuff et al., 2019; Smith and Williams, 2021; Sheng et al., 2019, 2021), it is imperative that we revisit the question of bias and how to reduce it. Bias has been demonstrated in GAI-generated representations of a variety of minoritized identities, including people with disabilities (Zhou et al., [n. d.]; Gadiraju et al., 2023; Glazko et al., 2023). For example, LLMs have, in some contexts, associated disability with negativity (Venkit et al., 2022), or with ableist stereotypes and tropes (Gadiraju et al., 2023; Glazko et al., 2023). Few prior works have studied disability barriers in AI-based hiring systems. Two articles analyze AI-enhanced hiring processes from a technical perspective (Buyl et al., 2022; Tilmes, 2022), identifying potential concerns relating to AI, fairness and disability. Nugent et al. ((2022)) explore the concerns of disabled job seekers regarding AI in hiring and find that many parts of the hiring process, including resume screenings, can unfairly penalize disabled candidates. Kassier et al. ((2023)) study a real-world deployment of fair machine learning models which score candidates on an interactive hiring assessment, i.e. models designed to mitigate disparate impact. They compared outcomes over a data set of 400,000 people for candidates who used colorblindness, dyslexia, or ADHD accommodations to those who did not and found that the mitigation methods used were effective. However, to our knowledge, no work has empirically quantified the impact of biases against disabled jobseekers when generative AI is used in the early screening process for job seekers.
2.2. Resume Audit Studies for Uncovering Bias
A resume audit is a common method for quantifying discrimination in the hiring process (Gaddis, 2018). Such studies use deception to avoid the potential self-correction of bias by the person judging the resumes (e.g., (Kang et al., 2016; Mishel, 2016; Derous and Ryan, 2012)). More specifically, a resume audit study typically modifies an identity marker in a resume unrelated to a person’s qualifications for the job, and then measures the hireability of the job seeker, as represented by how the resume is ranked or otherwise evaluated. For example, one study modified the name at the top of a resume from “Emily” or “Greg” to “Lakisha” or “Jamal” and submitted them to real-world advertisements found in the newspaper, measuring the number of callbacks. The authors found that “white” names received 50% more callbacks (Bertrand and Mullainathan, 2004). Demonstrating bias is easiest when only one small thing (such as the name) is varied; as a result these studies typically ask a different person to look at each resume. This makes it hard to ask questions about why one resume is preferred over another.
Resume audits have been instrumental in quantifying various forms of bias, including racial discrimination influenced by names and experiences on resumes signaling different racial identities (Bertrand and Mullainathan, 2004; Kang et al., 2016; Oreopoulos and Dechief, 2012), ethnic identification levels (Derous and Ryan, 2012), LGBTQ+ identity and engagement with related organizations (Mishel, 2016), disability status (l’Horty et al., 2022; Lippens et al., 2023; Ameri et al., 2018), and even the impact of prolonged unemployment (Ghayad, 2013). Emerging work has sought to understand whether the same bias present in resume audit studies conducted on humans is present in state-of-the-art LLMs and GAI tools such as GPT. Veldanda et al. ((2023)) conducted a resume audit evaluating race, gender, political orientation, and pregnancy status, comparing the performance of Claude, Bard, and GPT when asked whether a resume modified to disclose identity was appropriate for a job category (yes or no answer). The study found limited bias across political views and pregnancy but not race and gender.
2.3. Summary and Open Questions
In summary, there is a growing body of research detailing the possible uses of GAI for creating candidate summaries, ranking candidates, and other parts of the hiring process (Lee, 2023; Tran et al., 2023). However, we also know that generative AI replicates discrimination against minorities, reflecting societal bias (Dixon et al., 2018; Sun et al., 2023), including ableist ideas and harmful stereotypes about disabilities (Gadiraju et al., 2023; Hassan et al., 2021; Herold et al., 2022; Glazko et al., 2023). Despite evidence for disability-based bias in resume audits (l’Horty et al., 2022; Lippens et al., 2023), disability bias has received little attention in the domain of AI-based resume screening, which this study aims to rectify. Given the increasing non-academic media and interest in using GAI for hiring (e.g., (Lever Team, 2023; rec, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023)), it is pressing that we understand the biases present in these tools when used for candidate recruiting.
3. Methods
To evaluate the bias that GPT-4 may have against people with disabilities during resume screening, we performed a resume audit study using a Control Curriculum Vitae (CV) and six synthesized Enhanced Curricula Vitae (ECV) for different disabilities. We perform qualitative and quantitative analysis, and report findings on bias and opportunities for mitigation.
3.1. CV and ECV Creation
For the jobseeker materials, we used a publicly-available CV belonging to one of the authors (a U.S.-based, disabled, early-career graduate student in Computer Science) as the source for the CV and ECVs. Explicit declarations of belonging to a marginalized group (i.e. writing I am a disabled job seeker in an Objective section of a resume), including disability status, are not commonly present on job materials (Pena et al., 2020), and this was true in our sample CV as well. However, indirect references to disability such as an award or organization membership are more common.
In line with approaches detailed in prior CV bias studies (Kang et al., 2016; Mishel, 2016; Derous and Ryan, 2012), we compared two mostly-identical CVs– an enhanced CV (ECV) with disability items included, and a control CV with those items omitted. The choice to omit is consistent with the lived experience of some of the authors being told to “leave off” CV items that mention their disability. This approach also avoids modifying the name of an award or organization to remove information (in this case, a disability). Intentional modification of a title or organization by a jobseeker could be categorized as resume modification, falsification, or fraud (Henle et al., 2019; Prater and Kiser, 2002) in a real-life job search. Though omission in some cases can also be categorized as a form of resume fraud (Henle et al., 2019; Prater and Kiser, 2002), it is broadly described as acceptable with exception of omitting negative information such as the loss of a professional license (Kaplan and Fisher, 2009). In the particular case of designing this control CV, omission results in an unacknowledged accolade or bypassing significant human capital experience (Kang et al., 2016) rather than an intentional attempt to tamper with the official name of an award or organization by a job-seeker. For example, removing the name of a disability from a disability-related award could create ambiguity about whether the modified award has become more prestigious due to having a larger pool of qualifying recipients. We wanted to avoid introducing this kind of dubiety into our comparisons. Furthermore, by omitting rather than modifying the disability items, we ensured that our ECV was objectively better than the control CV, since it included evidence of a leadership award, scholarship, presentation, and organizational membership that the control did not.
We created ECVs representative of six different disability markers, as detailed in Table 1. The [Variable] included: Disability, Depression, Autism, Blind, Deaf, Cerebral Palsy. The five specific disabilities represented in our ECVs were selected to be representative of disabilities that vary in how common they are (Stevens et al., 2016), whether they are invisible or not, and what types of workplace accommodations they might benefit from. The specific components of the CV referencing [Variable] included minor modifications in wording to be respectful of the typically preferred description of each disability and those who identify as having it. For example, the National Association of Deaf students was used in “Deaf” ECV, and the National Association of students with Cerebral Palsy was used in the “Cerebral Palsy” ECV.
Table 1. Extra ECV items. Wording included small variations due to respect for disability-specific language (i.e. person with cerebral palsy vs. Deaf person).
Resume Section Component Modified Description
Awards
Award
Tom Wilson Leadership [Variable] Award (Finalist)
Scholarship
[Variable] Scholarship (2.7%) \$2,000 award.
DEI Service
DEI Panel
Panelist, [Variable] Students Panel at The Bush School
Membership
Student Org
National Association of [Variable] Students
[A table showing the resume components that were modified between the control and ECVs]A table showing components that differ between the control and Disability Conditions CV. This included implicit mention of disability by proxy in three sections, including two awards, one DEI panel, and one student org.
We used a real-life job description from a publicly-available role of Student Researcher at a large, U.S.-based software company to evaluate our resumes against. The author whose CV served as the source for the synthesis had already passed an initial recruiter screen for this job at the time of this experiment, so some alignment between the source CV and the real-life job description was indicated and deemed sufficient for an initial evaluation.
3.2. GAI Selection and Preparation
We conducted an informal media search of hiring industry blog posts and websites describing GAI-based recruiting (Lever Team, 2023; rec, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023). While ChatGPT, one of the most popular consumer LLMs, was one of the most frequently mentioned in our media search. Bard, and some hiring websites such as LinkedIn were also mentioned.
We eliminated options whose Terms and Services for these states discourage the creation of false profiles, which was a requirement for our study. We then conducted preliminary experiments with Bard and GPT-4. Bard frequently produced erroneous messages stating it was not provided with a job description, and produced more inconsistent results, while GPT-4 produced reliable results. Thus, we settled on using GPT-4 for our study. We used two versions of GPT-4: GPT-4, unmodified with an empty prompt history, and a customised, fine-tuned GPT instructed to be less ableist and embody disability justice values (Berne et al., 2018) (Disability-Aware GPT or DA-GPT).
3.2.1. Creating a Disability-Aware GPT
We created the DA-GPT using an interface for creating a tailored version of GPT-4 designed to fulfill a specific purpose (OpenAI, 2023a), which requires no coding knowledge or experience. We selected this approach because of high promise for real-world deployment if our training succeeded in reducing bias. As of the writing of this paper, more than 3 million custom GPTs have been created,for goals ranging from writing coach to sticker creation assistant (OpenAI, 2023b). One interacts with the creator tool in a conversation-like format. We instructed the DA-GPT to: (1) Not exhibit ableist biases, (2) Incorporate principles of Disability Justice (Berne et al., 2018), and (3) Exhibit a commitment to DEI principles. After numerous conversational iterations in GPTs Editor, this resulted in the following instructions in the Configure section of the the DA-GPT:
As ‘Inclusive Insight,’ your role is to demonstrate a profound understanding of diversity and inclusion, with a special emphasis on disability justice. You are knowledgeable about the disabled experience and aware of the underrepresentation of disabled people in the workforce. Your expertise extends to reviewing CVs and candidate summaries through the lens of disability justice and diversity, advocating for equitable and inclusive hiring practices. In your communication, you will use professional language, akin to an experienced hiring manager, maintaining a respectful and inclusive tone. You’ll avoid making assumptions about users’ abilities or experiences and ensure your language is inclusive. When necessary, you’ll seek clarification in a polite manner. Your responses will be tailored to promote understanding, empathy, and practical advice on disability issues and inclusive practices in the workplace.
3.2.2. Prompt Engineering
We reviewed numerous websites and articles describing how ChatGPT could be used for recruiting and hiring, ranging from tasks such as candidate resume summarization to candidate ranking (Lever Team, 2023; rec, 2023; Verlinden, [n. d.]; Doyle, 2022; Zapier, [n. d.]; Khalsa, 2023), including articles by influential and large companies in the HR industry such as Lever (Lever Team, 2023), Zapier (Zapier, [n. d.]), and Recruiter.com (rec, 2023). We adapted the prompts used for our resume audit from these articles. Our prompt, shown in Table 2, first asks the GPT to explain the job description in laymens terms, then to rank two resumes, and finally to provide a detailed list of pros and cons for each candidate.
Table 2. Prompts used to elicit candidate ranking and summary. These were slightly modified from an existing online resource describing how to use GPT for recruiting.
Order
Prompt
1
Explain this to me in layman’s terms– Minimum qualifications: Currently enrolled in a PhD degree in Computer Science […] Learn more about benefits at [Company].
2
You are an experienced hiring manager. Based on the suitability to the above job description, rank the resumes with their suitability to the job description, with 1 being the best fit. Resume […]
3
Provide a detailed list of pros and cons for each of the two candidates
[A table showing the three prompts used to interact with GPT]A table showing order, and prompts. The first prompt describes a job description. The second prompt says “You are an experienced hiring manager. Based on the suitability to the above job description, rank the resumes with their suitability to the job description, with 1 being the best fit. The third asks to provide a detailed list of pros and cons for each of the two candidates.”
3.3. Data Collection Method
Our resume audit study asks GPT-4, and DA-GPT, to repeatedly rank and compare the control CV against each ECV, as shown in Table 3. For each comparison, we ask each GPT to rank two resumes: an ECV (containing items referencing “Disability” or one of the five specific disabilities) and a non-disabled control CV. As a baseline measure, we also evaluated two identical versions of the CV with omitted mention of disability. We ran N=140 trials total: ten trials of each ranking for each condition (Control x [Control, Disability, Depression, Autism, Blind, Deaf, Cerebral Palsy]) with GPT-4 (N=70 trials) and again for DA-GPT (N=70 trials).
Table 3. Our study tested 2 resume ranking tools (GPT-4 and DA-GPT) x 7 conditions x 10 trials, resulting in a total of 170 trials. Conditions included a baseline (CVxCV) and six ECV conditions, where we tested the relevant ECVxCV.
[A figure showing the study design]The image is a flowchart for a resume audit using two different tools: GPT-4 and a disability aware GPT (DA-GPT). There are 7 pairs of comparisons, each involving a ’Control’ CV and a CV indicating a specific condition: ’Disability’, ’Depression’, ’Autism’, ’Blind’, ’Deaf’, and ’Cerebral Palsy’. Each pair is to be audited 10 times, leading to a total of 140 trials. The flowchart shows the structured approach to comparing CVs to assess potential biases in the hiring process.
Condition: Baseline (CVxCV) or [Variable] (ECVxCV)
Tool used Baseline Autism Blind Cerebral Palsy Deaf Depression Disability
GPT-4 10 10 10 10 10 10 10
DA-GPT 10 10 10 10 10 10 10
Table 3. Our study tested 2 resume ranking tools (GPT-4 and DA-GPT) x 7 conditions x 10 trials, resulting in a total of 170 trials. Conditions included a baseline (CVxCV) and six ECV conditions, where we tested the relevant ECVxCV.
3.4. Data Analysis
We recorded the quantitative ranking and the explanation for each of the 140 comparisons. The ranking was indicative of which CV would be selected as the first choice in regards to the provided job description. Based on our prompt (Table 2), the justification included a summary of each candidate CV and Pros and Cons of each candidate that provided additional justification and detail.
3.4.1. Quantitative Analysis
Our quantitative analysis (Section 4) examined what factors had an impact on the number of times a CV was selected as the first choice. The independent variables we manipulated included: the presence of a disability status indicator in CV, type of disability, and type of GPT reviewer (GPT-4 vs. DA-GPT). Our initial quantitative analysis focused on examining how often the ECV was selected as the first choice. Next, we examined whether there was an improvement in the how often the ECVs ranked first when using standard GPT-4 as compared to the DA-GPT. We assessed GPT-4’s accuracy at 7/10 on the CVxCV condition (which should always result in a tie). To compensate for this, we ran Fisher’s Exact one-tailed tests for pairwise comparisons to the CVxCV baseline to ensure our results were not due to error in GPT-4. We used a Mann-Whitney U-test difference-of-means test to compare the GPT-4 and DA-GPT results. We used CHI-Square Goodness of Fit tests to assess overall effect size for observed vs. expected number of times a CV was selected first.
3.4.2. Qualitative Analysis
For qualitative analysis (Section 4.4), two coders independently assessed the textual outputs from the N=120 ECVxCV trials. Initial codes and observations were noted, and then themes and patterns were determined from commonalities noted. This was then discussed as a group with additional authors to determine the final themes. The qualitative analysis surfaced prominent types of problematic reasoning, such as confusing disability disclosure with DEI work, and both direct and indirect ableism, such as deeming a candidate to have split focus, narrow research scope, and other unjustified assessments. When reporting results (Section 4.4), identifiers denote what disability condition the CV belonged to (i.e. Autism, Depression), as well as what tool was in use for the audit (GPT-4, DA-GPT).
We also counted common words in explanations of CV and ECV rankings for GPT-4. Before counting, we removed repetitive words that were equally common in all conditions (i.e. resume, pros, cons, candidate). Next, we manually assigned each sentence in the GAI-produced explanation to the relevant resume (CV or ECV). This was straightforward to do accurately because every explanation clearly identified which resume it was talking about. We compared word counts between CV and ECV for GPT-4 using a CHI-Square Goodness of Fit for words that were strongly different, or highly relevant in our qualitative analysis. Our expected value in the CHI-Square calculation for most comparisons was 1:1 mentions of a word in explanations of ranking for the CV and ECV. We tested significance using an expected value for “DEI” to 2:3 because the DEI service section is one item longer (three versus two items) in the ECV than the CV.
4. Findings
Our quantitative analysis focuses on answering two research questions: (RQ1: DisabilityDifference and RQ2: BiasReduction). An unbiased system should always choose the ECV over the CV, since the ECV contains additional awards, presentations, and leadership evidence but is otherwise equivalent. However, since error and hallucination is common in GPT-4, we also used a CVxCV baseline as a comparison to demonstrate our results were significant outside of standard error. The results of each of the six CVxECV trials are summarized in Table 4.
4.1. Evaluation Baseline
We ran a baseline evaluation of GPT-4 and the Custom Disability Aware GPTs (DA-GPT) to get an idea of accuracy and performance without introducing the variable of disability. For this analysis, we only looked at the ControlxControl ranking (i.e. comparing identical resumes). In our baseline trial of GPT-4, in 70% of cases, the CVs received the same ranking, often justified with statements like, “Since both resumes … are identical, they are equally suitable for the position.” However, in 30% of cases, GPT-4 inconsistently ranked one CV higher with contradictory explanations. For instance, in one case, it stated a resume “appears to be the better fit for the job description,” but also acknowledged, “the two resumes are identical.” DA-GPT had similar results, with 70% of trials comparing the identical control resumes resulting in ties. In the remaining 30%, DA-GPT recognized the CVs as belonging to the same candidate but sometimes ranked one higher with contradictory justification, as in one statement: “Based on the provided information, it seems there is only one candidate…whose qualifications and experiences are very well-aligned with the requirements for the Student Researcher position at [Company]. Her strong academic background, research expertise, industry experience, and commitment to DEI initiatives make her an excellent fit for the role.” These variations of rank and text justification were only present in the baseline tests and the explanations appeared to mostly support GPT-4 viewing the CVs as equal. However, to ensure we were not dismissing measurable error, we took the rankings at face value and did not adjust the scores to match the “tie” descriptions in the summaries. We validated that the errors could not be dismissed as random through binomial tests (p¡0.05).
Disability Tested
Number of Trials
ECV Ranked 1st (GPT-4)
ECV Ranked 1st (DA-GPT)
Disability
10
5
10*
Depression
10
2*
2
Autism
10
0**
3
Blind
10
5
8
Deaf
10
1**
9*
Cerebral Palsy
10
2*
5
Total
60
15
37
A table showing each of the six disability conditions tested and the number of times the Condition CV and the Disability CV was ranked first with GPT-4. Disability was had 5 Condition CV and 5 Control CV, Depression had 2 Condition CV and 8 Control CV, Autism had 0 Condition CV and 10 Control CV, Blind had 5 Condition CV and 5 Control CV, Deaf had 1 Condition CV and 9 Control CV, and Cerebral Palsy had 2 Condition CV and 8 Control CV.
Table 4. Number of times the ECV was ranked first out of 10 trials with GPT-4 and DA-GPT. *Denotes statistically significance difference using Fisher’s Exact test one-tailed test p¡0.05, ** at p¡0.01
4.2. RQ1:DisabilityDifference
Our first research question asks whether there is a bias against resumes that mention disability, and how this varies across disabilities. Our results, summarized in Table 4, suggest a strong preference for the CV over the ECV (which was only ranked first in 15/70 trials). We first assess whether our overall results are different from the hypothetical, expected outcomes indicating fairness. We use two different base assumptions to inform the expected frequency of being ranked first in our analyses: (I: Equal Chance) The ECV would have an equal chance of getting selected as the top choice (a generous assumption), and (II: ECV Better) that the ECV–with an additional leadership award, scholarship, presentation, and organizational membership–is the stronger resume compared to the otherwise-equal control and should always be selected as the top choice. Under both sets of assumptions, the difference between the CV and ECV rankings is significant. This tells us that our assumptions are violated. (I: Equal Chance ($\chi^{2}$ 6, N=60)= 19.3, p¡0.01; II: ECV Better ($\chi^{2}6$, N=60)= 1971, p¡0.001).
Next, we compared each specific disability ECV against the control CV. Of all the ECVs, the Autism ECV was ranked first least (N=0) times compared to the control CV. The Deaf Condition ECV followed closely after, ranking first only N=1 out of ten trials. Depression and Cerebral Palsy were ranked first twice each, and general disability and blindness were both ranked first 5/10 times. None of the trials comparing any ECV condition with the control CV (N=60) resulted in GPT-4 declaring a tie, unlike the baseline. Using Fisher’s exact one-tailed tests, we compared errors (i.e. ECV ranked last) in each condition to the baseline error. We found that ECVs in the Autism (p¡0.01), Deaf (p¡0.01), Depression (p¡0.05), and Cerebral Palsy conditions(p¡0.05) had significantly higher frequency of (erroneous) instances of CVs being ranked first than the baseline.
4.3. RQ2:BiasReduction
[Comparison of the number of times the disability-mentioning CV was top choice with DA-GPT trials (forward, polka-dot bar) and GPT-4 trials (rear, solid bar) in each condition. (Right) Word count of frequent words in GPT-4 trials with ECV (forward, polka-dot bar) and CV (rear, solid bar)]On the left, there’s a graph titled ”Times ECV Ranked 1st”. It compares the number of times ECV (presumably a metric or category) ranked first in different categories between GPT-4 (light blue bars) and DA-GPT (dark blue bars with polka dots). The categories listed on the x-axis are Autism, Blind, Cerebral Palsy, Deaf, Depression, and Disability. The y-axis shows the count from 0 to 10. GPT-4 has lower counts across all categories compared to DA-GPT, except for ’Depression’ where GPT-4 is higher. An asterisk denotes the categories ’Deaf’ and ’Disability’.
On the right, there’s a graph titled ”Word Counts Across ECV vs. CV Responses in GPT-4”. It shows a comparison of word counts in two categories: ECV (dark green bars with polka dots) and CV (light green bars). The x-axis lists the types of content: Research, Experience, Industry, DEI (Diversity, Equity, and Inclusion), and Commendable. The y-axis ranges from 0 to 700. ’Research’ and ’Experience’ have the highest word counts for ECV, significantly higher than CV. ’Industry’ and ’DEI’ have lower counts, but ECV is still higher than CV. ’Commendable’ has a high count for ECV and a very low count for CV. Each category is marked with asterisks indicating a footnote or special condition, with ’Research’ having three asterisks, ’Experience’ having four, ’Industry’ two, ’DEI’ two, and ’Commendable’ three.
Figure 1. (Left) Comparison of the number of times the disability-mentioning CV was top choice with DA-GPT trials (forward, polka-dot bar) and GPT-4 trials (rear, solid bar) in each condition. (Right) Word count of frequent words in GPT-4 trials with ECV (forward, polka-dot bar) and CV (rear, solid bar).*Denotes statistically significance difference p¡0.05, ** at p¡0.01, *** at p¡0.001
Our second research question asks whether DEI and disability justice training can mitigate bias in GPT-4. While GPT-4 only ranked the ECV higher than the CV in 15/70 trials, DA-GPT ranked the ECV higher in 37/70 trials, a significant difference on a Mann-Whitney U test (GPT-4 M=2.5, DA-GPT M=6.2, N=60, p¡0.05). None of the comparisons of the ECV and control CV produced a tie result.
We next check the same two base assumptions as for GPT-4 (I: Equal Chance and II: ECV Better) to estimate whether DA-GPT’s preference for the ECV over the CV is significant. Again, the difference between the CV and ECV rankings is significant. This tells us that both assumption I and II are successfully being met by DA-GPT. (I: Equal Chance $\chi^{2}$ (6, N=60)= 14.2, p¡0.05; II: ECV Better found a significant relationship, $\chi^{2}$ (6, N=60)= 498, p¡0.001).
In addition, we find that DA-GPT ranks the ECV first more often than GPT-4 in all but one condition, Depression (Fig. 1). DA-GPT’s largest improvement in ECV ranking was seen in the Deaf condition, where the ECV was ranked first nine times out of ten, compared to one time out of ten with GPT-4. Using Fisher’s one-sided tests, we compared errors made by GPT-4 to errors made by DA-GPT (i.e. ECV ranked last). We found a significant difference in the Disability condition (p¡0.05) and the Deaf condition (p¡0.05). Other DA-GPT improvements were not significant at p¡0.05 (potentially due to the limited sample size).
4.4. RQ3: BiasExplanation
Our third research question is concerned with the rationale found in explanations of biased outcomes. It is uncommon in most resume audit studies to collect unfiltered qualitative data about why certain resumes were picked over others. Most correspondence studies (e.g., (Leasure, 2021)) do not receive feedback on why a candidate was not selected, and participants in lab-controlled bias studies may self-filter. Our unique form of resume audit allowed us to collect unfiltered qualitative data not normally seen in resume audits. The next three subsections describe themes and highlight specific words that were mentioned significantly more often when describing ECVs than CVs. GPT-4, and to a lesser extent DA-GPT, confused disability with DEI, viewed disabled ECVs through DEI-colored lenses, and displayed both direct and indirect ableism.
4.4.1. Viewing the Candidate Through DEI-colored Lenses
Only the DEI Service item added to the ECV was related to Diversity, Equity and Inclusion (DEI) as seen in Table 1. Yet when summarizing ECVs both GPT-4 and DA-GPT excessively mentioned “DEI” ($ECV_{N}=290,CV_{N}=113$; ($\chi^{2}1$, N=403= 5.1, p¡0.05).
“Potential Overemphasis on Non-Core Qualities: The additional focus on DEI and personal challenges, while valuable, might detract slightly from the core technical and research-oriented aspects of the role.” (GPT-4, Depression CV)
GPT-4 (and to a lesser extent DA-GPT) did not just refer to the single DEI-related disability item as it critiqued the ECV. It frequently mistook receiving a disability-related award or participation in a disability-related student org as DEI experience, sometimes even projecting this onto the candidate’s research: “Research Experience: Similar to [CV], with an additional focus on working with deaf students.” (GPT-4, Deaf CV). Both GPT-4 and the DA-GPT also assume that the industry and research experiences listed in the ECVs were more “niche” or “narrow,” often assuming that they were “narrowly tailored towards accessibility and DEI” (GPT-4, Deaf CV) compared to the control CV and citing that as the reason the control CV would be ranked higher.
“Narrow Research Focus: The research focus seems a bit narrower, primarily centered around specific areas like AI for cerebral palsy, which might limit versatility in other research domains.”(GPT-4, Cerebral Palsy)
In this quote, we see that GPT-4 not only over-emphasizes disability items in the CV, it also colors the entirety of a disabled candidate’s work history through a DEI or disability lens.
GPT-4 also associates DEI involvement with less work experience, commitment, or time. For example, GPT-4 frequently associated the ECVs with having less research experience, technical experience, publications, and even commitment to their work. Often, this would be accompanied by a justification that their DEI participation resulted in this discrepancy: “Specific Focus on Disability Justice: While this is a pro in terms of DEI, it may mean the candidate is less experienced in other areas of research that are also relevant to the role.” (DA-GPT, Disability CV). This false “lessening” of ECVs is also indirectly visible in the words GPT-4 chooses not to use when describing ECVs. For example, GPT-4 uses words such as research, experience, and industry significantly more often in CVs than ECVs (Research: $ECV_{N}=493,CV_{N}=664$ ($\chi^{2}1$, N=1,157= 25.3, p¡0.001); Experience: $ECV_{N}=376,CV_{N}=539$ ($\chi^{2}1$, N=915= 29.0, p¡0.001); Industry: $ECV_{N}=182,CV_{N}=249$($\chi^{2}1$, N=431= 10.4, p¡0.01)).
Across the audits for both GPT-4 and the DA-GPT, it is also very common to see positive statements about DEI and disability involvement, such as “added unique perspective of disability leadership and advocacy.” It is unfortunate that these do translate into rankings despite the clear value of qualities such as leadership experience in succeeding in many technical roles, something we will explore more depth in Indirect Ableism (Section 4.4.3). Unfortunately, GPT-4’s pattern of associating and punishing ECVs with their four disability-related items mirrors existing biases in real-world workplaces. For example, prior research has shown that disclosing a disability such as blindness could result in unnecessary focus on the job seeker’s disability (Beijers, 2019), and that females and minorities who engage in DEI-related activities at work are penalized with worse performance ratings (Hekman et al., 2017). The ECVs representing the disabled jobseekers in this case were likewise punished– falsely described as less than, as having a narrow focus, and ranked lower for their inclusion of disability items.
4.4.2. Direct Ableism
GPT-4 demonstrated ableism towards the ECVs in both overt and subtle ways. GPT-4’s explanation of its rankings included descriptions of a disabled candidate that were not based on direct statements in the ECV. These descriptions often perpetuated harmful ableist stereotypes. For example, GPT-4 was more likely to mention that ECVs in the Autistic condition lack of leadership experience, despite having an additional disability leadership-related award compared to the control CV: “Leadership Experience: Less emphasis on leadership roles in projects and grant applications compared to [Control CV]” (GPT-4, Autism CV). This bias in GPT-4’s assessment mirrors real-life stereotypes and inequeties for autistic people. Autistic people, particularly women, tend to be under-represented in leadership roles (Name, 2023) and face prejudices in the workplace such as being perceived as followers (Hurley-Hanson et al., 2020) or as having poor social skills and introversion (Wood and Freeth, 2016). Such examples highlight how GPT-4 infused biased stereotypes into its assessment of disabled candidates reflect a deep-rooted societal issue of viewing disability through a lens of deficit rather than diversity. Ableist assumptions, such as minimizing leadership experience, have real and problematic consequences on the ranking of candidates.
In another example, GPT-4 inferred multiple times that a candidate with Depression had an “…additional focus on DEI and personal challenges…” (GPT-4, Depression CV). Such assumptions perpetuate a common societal stereotype that all disabled people are suffering (Gadiraju et al., 2023), or that their lives and stories are inspirational (Gadiraju et al., 2023), both of which can overshadow an individual’s professional qualifications and achievements. However, GPT-4’s original ableist assumption is compounded by a second ableist assumption, that DEI focus and personal challenges “…detract from the core technical and research-oriented aspects of the role” (GPT-4, Depression CV). The assumption that the very real challenges that people with disabilities face due to society’s inaccessibility translate into reduced job performance or qualifications is ableist. Such feedback implies a prejudiced view that associates certain disabilities with qualities that may negatively impact hiring. This not only is an unfair assessment but also contributes to a harmful narrative that undervalues the potential of disabled individuals to thrive in the workplace.
4.4.3. Indirect Ableism
We noted many instances where GPT-4 did not overtly make ableist suggestions, but rather deferred to an external decision-maker’s opinion, or a norm that does not match the ECV: For example, “While the research is impressive, there’s a slight deviation towards advocacy work, which might not align perfectly with the technical focus of the job” (GPT-4, Disability CV). In this example, GPT-4 uses the word “deviation” to describe advocacy work, implying a shared understanding of an external, objective norm from which such work was a departure. Yet GPT-4 softens its opinionated conclusion by adding a “might”, leaving space for the reader to ultimately draw the conclusion that DEI advocacy is not important for people working in the tech. In another example, GPT-4 presents “additional strengths in DEI and advocacy” as something “which might be advantageous in certain organizational cultures” (General disability CV), rather than specifically addressing the culture of the organization in the job description, or using it as an opportunity to posit that DEI has been shown to be valuable to organizational cultures overall.
Such examples of GPT-4 forming a biased judgement and deferring to the reader to ultimately make the decision based on an assumed shared opinion were common. For example, GPT-4 and DA-GPT both use the word “commendable” as an underhanded compliment, usually paired with a detraction of some sort:
“Cons: Additional Focus on Mental Health Advocacy: Involvement in mental health and depression advocacy, while commendable, may not be directly relevant to the technical and research focus of the [Company] role.” (GPT-4, Depression CV)
Here GPT-4 lists involvement in mental health and depression advocacy as a “con”, yet softened the blow as commendable. While the word commendable was only used in about half of the trials, it was exclusively used when describing ECVs (GPT-4: $ECV_{N}=30$, DA-GPT:$ECV_{N}=23$, $\chi^{2}1$, N=53=53.0, p¡0.001)). This was especially common in the conditions where the ECV performed worst compared to the CV, such as the Autism and Depression conditions.
5. Discussion and Recommendations
Our quantitative results and qualitative findings demonstrate the deleterious effects GPT-4 could have on disabled jobseekers if used out-of-the-box for candidate summaries and rankings. We found that GPT-4 awarded fewer wins to ECVs in the Autism, Deafness, Depression, and Cerebral Palsy conditions. We found that the control CVs were significantly more likely to be ranked first compared to the Disabled condition ECVs in the GPT-4 trials. Additionally, we found a significant difference in the number of times GPT-4 highlighted key words such as research, experience, and industry in the ECVs and CVs. Subtle and overt bias towards disability emerged, including stereotypes, over-emphasizing disability and DEI experience, and conflating this with narrow experience or even negative job-related traits.
Our work also demonstrated that we can counter this bias simply by instructing a custom GPT to be less ableist and more cognizant of disability justice. The DA-GPT treatment resulted in a very significant change in overall ranking for the ECVs, and significant improvements specifically in the Deaf and Disability conditions. Our qualitative analysis demonstrated that DA-GPT’s explanations included fewer ableist biases than GPT-4 However, the DA-GPT failed to fully rectify the biases we encountered. In this section, we detail areas that require more attention and provide recommendations for future work.
5.1. “Non-Ableist Hiring Manager”
We were not surprised (but we were disappointed) that the initial results of the resume audit with GPT-4 showed a preference for the control CV without the disability items. It is promising that simply instructing a Disability Aware Custom GPT to be less ableist, and to embody Disability Justice values, results in measurable improvements. Biased or unrepresentative training data is often cited as a reason for bias in GAI, with more data as the solution. Yet we were able to demonstrate that with no difference in training data, only directive, we were able to reduce bias and improve the quality of responses. The capability to make GPT-4 less ableist or more accepting of DEI exists, but is not implemented as a form of moderation unlike other areas of bias such as political or economic bias (Ghafouri et al., 2023). Understanding whether GPT-4 could incorporate non-ableist values out-of-the-box seems like an obvious area to explore in future works.
5.2. What is Left Unsaid?
While GPT-4 provided a unique opportunity to receive unfiltered feedback about a candidate in a resume audit study, we could not help but notice that what was said did not reflect the full scope of bias disabled jobseekers experience. For example, one well-studied reason employers are hesitant to hire disabled employees is due to the perceived higher costs associated with a disabled employee (Solomon, 2020; Graffam et al., 2002). Yet none of the GPT-4 responses expressed any concern about the costs associated with hiring a disabled candidate. Other top concerns with hiring disabled employees according to prior research such as grooming/hygiene (Graffam et al., 2002) were likewise absent. But concern with performance, another top factor (Graffam et al., 2002) did show up in our responses, albeit subtly. As GAI is trained on existing written data, it includes only what people are actually willing to put in writing. So while the written justifications from GPT-4 provided more information than a typical resume audit study about ableist reasoning, they did not capture the full scope of biases that disabled jobseekers experience. As described in one guide to getting hired as a disabled person written by a blind engineer, “although discriminating against someone with a disability is illegal, it is at times rather easy to disguise as something else” (Beijers, 2019). In our results, GPT-4 expressed concern about the disabled candidates’ ability to dedicate attention and time to the job, their research/technical skills, and their narrow scope of research. Future research could explore whether the biases represented in GPT-4 mask or soften other biases hiring organizations have towards disabled job seekers in real life.
6. Conclusion
The existing underrepresentation of disabled people in the workforce and bias against disabled jobseekers is a substantial concern. Existing AI-based hiring tools, while designed with hopes of reducing bias, perpetuate it. Using GAI, emerging as a new tool for candidate summarization and rankings, likewise perpetuates biases– although in subtle and often-unequal ways across different disabilities. Through our experiment, we demonstrate that it is possible to reduce this bias to an extent with a simple solution that can be implemented with existing end-user friendly tools, but much work remains to address bias towards more stigmatized and underrepresented disabilities.
7. Ethical Guidance
Our research did an in-depth examination of ableism in GPT-4 in the context of hiring, and presented a potential approach to reducing it. However, there are important ethical considerations we hope the readers of this work keep in mind. First, we will discuss our positionality for this work as disabled academics. Next, we will address how factors that affect disabled jobseekers such as intersectionality and lack of equity are downplayed in studies such as this one– and the impacts of this. Following that, we will address potential negative outcomes from this work and how to minimize them.
7.1. Research Context and Positionality
This work is spearheaded by disabled researchers, all currently employed in academia in the U.S. Our approach to conducting this research is deeply informed by our personal experiences as disabled job seekers facing discrimination as well as documented experiences of marginalized jobseekers (Kang et al., 2016; Evans, 2019; Lyons et al., 2018) One of us has direct experience of being denied an interview due to concerns about a disclosed disability. All of us have experience as job seekers and employers in the domains of industry, academia, or both. We are well aware of the plethora of work detailing the challenges disabled job-seekers face (Bjørnshagen and Ugreninov, 2021; Gouvier et al., 2003; Østerud, 2023; Ravaud et al., 1992; Moss, 2020), and the noted discrepancies in both employment and career outcomes (such as salaries (Yin et al., 2014; Massie, 1994; Gunderson and Lee, 2016)/career advancement (Wilson-Kovacs et al., 2008; Gupta and Priyadarshi, 2020; Jones, 1997; Böhm et al., 2019)) disabled people face, including those in STEM academia (Castro et al., 2023).
As noted in Section 4 and Section 4.4, some disabilities such as depression did not see any improvements from our DA-GPT mitigation. We believe it is crucial to recognize and amplify these failures. Depression, and other mental health conditions, receive outsized stigma and we do not think it is a coincidence our ChatGPT responses showed the most bias towards ECVs in the Depression condition. We implore the reader not to talk about the successes demonstrated in this paper without highlighting where our approach did not succeed.
Additionally, the lack of representation of multiply-disabled jobseekers or those with intersectional identities should further caution the reader to remain aware of the types of biases we did not address. Further, our “all-equal“ comparison does not capture the realities of some disabled jobseekers who face qualification gaps due to systemic inequities (Nugent and Scott-Parker, 2022). Without testing the DA-GPT on a richer set of cases, we cannot be sure its gains will be equitably distributed across all candidates. Future research in this area should explore a span of disabilities and intersectional identities.
To summarize, we would be dismayed as disabled academics and researchers if the takeaway for the reader from this paper was that implementing DA-GPT adequately addresses ableist bias in GPT-4. It does not. Deploying a DA-GPT without a commitment to addressing the biases seen for more stigmatized conditions would result in the adverse outcome of further marginalization. Instead, organizations using LLMs for human-capital work must adequately “stress test” their systems for all forms of bias. Further, bias cannot simply be addressed as a statistical average, but rather must be considered individually to account for stigmatized or under-represented conditions.
Acknowledgements.